OpenStudy (anonymous):

When 100 cm^3 of 2.0 mol dm^-3 hydrochloric acid is added to 100cm^3 of 2.0 mol dm^-3 aqueous sodium hydroxide, the temperature rise is 8.0 degree Celsius. Which of the following combinations produces the same temperature rise? (i) 50cm^3 of 1.0 mol dm^-3 HCl + 50 cm^3 of 1.0 mol dm^-3 NaOH (ii) 50 cm^3 of 1.0 mol dm^-3 HCl + 100 cm^3 of 1.0 mol dm^-3 NaOH (iii)100 cm^3 of 4.0 mol dm^3 HCl + 100 cm^3 0f 4.0 mol dm^-3 NaOH (iv) 200 cm^3 of 2.0 mol dm^-3 HCl + 200 cm^3 of 2.0 mol dm^-3 NaOH.

OpenStudy (anonymous):

Shouldn't we compare the number of moles here? I think none of the options is correct. >_< Is it so?

OpenStudy (mos1635):

almost calc given moles then determine in wich case they are the same moles reacting

OpenStudy (anonymous):

I did the same. But none of the options give the same number of moles. >_<

OpenStudy (mos1635):

HCL + NaOH -> NaCl + H2O given :0.02 0.02 reacts 0.02 case1:0.05 0.05 reacts 0.05 (regected) case2:0.05 0.01 reacts 0.01 (regected) check the oter two

OpenStudy (mos1635):

ooops!!!!!! that is not enough we mast consider volume of final solution!!!! so given: reacts 0.02 moles and heat rises 200cm^3 of sol to 8 degrees case1 0.05 100 10 got it??

OpenStudy (mos1635):

nice question!!!!!!!

OpenStudy (mos1635):

back in 2,345,346,637,663,856,672.34 feftosec

OpenStudy (anonymous):

Didn't get that. >_<

OpenStudy (anonymous):

@mos1635 : Can you please explain it again?

OpenStudy (mos1635):

HCL + NaOH -> NaCl + H2O given :0.02 0.02 reacts 0.02 final volume 200 case1: 0.05 0.05 reacts 0.05 final volume 150 case2: 0.05 0.01 reacts 0.01 final volume 200 case3: 0.04 0.04 reacts 0.04 final volume 200 case4: 0.04 0.04 reacts 0.04 final volume 400 so far so good??

OpenStudy (mos1635):

given : reacts 0.02 produses heat Q case1: reacts 0.05 produses heat Q1=2.5Q case2: reacts 0.01 produses heatQ2=Q/2 case3: reacts 0.04 produses heatQ3=2Q case4: reacts 0.04 produses heatQ4=2Q

OpenStudy (mos1635):

given : heat Q rises 200 cm^2 solution up 8 degres case1: heat Q1=2.5Q rises 150cm^2 solution up θ1=? degres case2: heatQ2=Q/2 rises 200 cm^2 solution up θ2=? degres case3: heatQ3=2Q rises 200 cm^2 solution up θ3=? degres case4: heat Q4=2Q rises 400cm^2 solution up θ3=? degres

OpenStudy (mos1635):

θ1=40/3 θ2=4 θ3=16 θ4=8

OpenStudy (mos1635):

all that assuming that densidy of final solutions aproximatly the same so Mi=ρ*Vi

OpenStudy (mos1635):

you can also use Qi=Ci*Mi*Δθi Qi=C*ρ*Vi*Δθi

OpenStudy (anonymous):

I'm sorry. But I have never used these formulas. But I understood the part upto the heat production.

OpenStudy (mos1635):

ok rise of temp is propotional to absorved heat and rise of temp is counter propotional to mass of solution (in our case qe use Volume)

OpenStudy (mos1635):

is that makes any sence?

OpenStudy (anonymous):

Umm. Ok! I think I got it. Thankyou so much! :D The final anwerd would be (iv) then?

OpenStudy (mos1635):

iv it is

OpenStudy (anonymous):

BTW, can you tell me what those formulas actually are? I mean what each letter denotes?

OpenStudy (mos1635):

Q= absorved heat (Joule) m= mass of body(Kg) c heat capasity (joule/Kg*K) Δθ =deferance in tempurature(K)

OpenStudy (anonymous):

Ok, I know this formula. But does that 'p' in the previous equation stand for density?

OpenStudy (mos1635):

yes